Which is a point that lies on the graph of the equation: 2x-y=8?
A (3,-2)
B (-2, 6)
C (-3,-2)
D (-2,-6)

Found 3 solutions by solver91311, Leaf W., Edwin McCravy:
Answer by solver91311(24713)



The first value in an ordered pair is the x-value. The second value is the y-value. Substitute the first value in each of the ordered pairs in place of x and substitute the second value in each of the ordered pairs in place of y. Do the arithmetic each time. If the result is a true statement, then the point is on the graph, otherwise it is not on the graph.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Leaf W.(135)

Knowing that a coordinate point follows the format (x,y), simply plug the values of x and y into the equation for each. If the equation is true, that is your answer.
A. (3,-2)
==> 2x - y = 8
==> 2(3) - (-2) = 8
==> 6 + 2 = 8
==> 8 = 8
This is true, so (3,-2) is a point on the equation 2x - y = 8, and your answer.
B. (-2,6)
==> 2x - y = 8
==> 2(-2) - 6 = 8
==> -4 - 6 = 8
==> -10 = 8
This is false, so (-2,6) is neither a point on the equation, nor your answer.
C. (-3,-2)
==> 2x - y = 8
==> 2(-3) - (-2) = 2
==> -6 + 2 = 2
==> -4 = 2
This is false, so (-3,-2) is neither a point on the equation, nor your answer.
D. (-2,-6)
==> 2x - y = 8
==> 2(-2) - (-6) = 8
==> -4 + 6 = 8
==> 2 = 8
This is false, so (-2,-6) is neither a point on the equation, nor your answer.
***Therefore, the correct answer is A. (3,-2)

Answer by Edwin McCravy(20055)

A (3,-2)
B (-2, 6)
C (-3,-2)
D (-2,-6)
We have to try them all out and see:

Let's try A (3,-2)

That means to substitute (3) for x and (-2) for y
in this equation and simplify:

     2x - y = 8
2(3) - (-2) = 8
      6 + 2 = 8
          8 = 8

That's true so A (3,-2) is the correct answer.
But let's see why the others are not the answer:

-----

Let's try B (-2,6)

That means to substitute (-2) for x and (6) for y
in this equation and simplify:

     2x - y = 8
2(-2) - (6) = 8
     -4 - 6 = 8
        -10 = 8

That's false so B (-2,6) is not the correct answer.

-----

Let's try C (-3,-2)

That means to substitute (-3) for x and (-2) for y
in this equation and simplify:

     2x - y = 8
2(-3) - (-2) = 8
     -6 + 2 = 8
         -4 = 8

That's false so B (-3,-2) is not the correct answer.

-----

Let's try D (-2,-6)

That means to substitute (-2) for x and (-6) for y
in this equation and simplify:

      2x - y = 8
2(-2) - (-6) = 8
      -4 + 6 = 8
           2 = 8

That's false so D (-2,-6) is not the correct answer.

Edwuin